# Predictions for Detecting a Turndown in the Baryonic Tully Fisher Relation

### v1

@@ 2.2 Spatially unresolved H i profiles
We use a fixed velocity resolution of km s-1 and measure each galaxy over the velocity range -200 to 200 km s-1. This velocity resolution is set to match the quality of Green Bank Telescope observations in Bradford et al. (2015). Our data cubes have fixed spatial axes spanning from -100 to 100 kpc across each galaxy with a resolution of 3.7 kpc, though the spatial resolution has little impact on the unresolved linewidth results. In this work, we do not consider beam smearing effects or observational noise.
@@ 2.3 Inclinations
At M M⊙, our simulated dwarfs are predominantly irregular rather than disky (Keith et al., 2025). For these lowest-mass dwarf galaxies, we have found that our orientation method does not always yield a result that is edge-on by visual inspection. It is possible that there is no orientation that yields an apparent edge-on result, given their irregular nature. Although the inclinations for these irregular galaxies are difficult to measure and correct for, the Gaussian H i profiles of these lower-mass dwarf galaxies do not vary much by viewing angle (as quantified in the errors presented in Section 3.2). Our inclination procedure with simulations allows us to include dwarf galaxies that would otherwise be removed from observational samples.
@@ 2.4 Definitions for Velocities and Radii
There are a number of different ways to measure rotation velocity, and at different radii. In this work, we aim to understand how the different methods impact the determined bTFR. The methods for velocities and corresponding radii in this work are:
@@
W and R: W, W, and W are the H i profile’s width at different percentages of the peak flux height; 10%, 20%, and 50%, respectively. The corresponding radius, R, is where the line width W intersects the circular rotation curve. For some cases, W may not exactly intersect the rotation curve (see Figure 2, top panel as an example). To obtain the equivalent radius, we calculate the difference between the rotation curve and W, and define R as the corresponding radius with the smallest difference within 5 kms-1 of W.
@@ 3.1 Comparing Stellar and H i Properties to Observations
After confirming that our masses are consistent with observations, we check the H i size-mass relationship in the bottom panel of Figure 4. The size is defined as the diameter where the H i surface density has dropped to 1 M⊙pc-2, similar to observations. We compare the H i size-mass relation from the simulated dwarf galaxies to the observations from Wang et al. (2016), which follows the trend from Broeils & Rhee (1997) (B97), and has been confirmed by recent WALLABY (Deg et al., 2024) and FEASTS results (Wang et al., 2025). Our galaxies fit in the 3- scatter of Wang et al. (2016), represented by the grey line and shaded region. Stevens et al. (2019) found that the H i size-mass relation is robust to gas stripping and galaxy population (centrals versus satellites), so matching this relation mostly confirms that our feedback is of appropriate strength and does not disrupt the H i disk.
@@
bTFR using the maximum halo speed: Modeling the halo’s circular velocity based on the total mass enclosed in spherical shells is common in bTFR observations (e.g., Trujillo-Gomez et al., 2018; McQuinn et al., 2022; Lelli et al., 2019), so we use this definition as our reference point. In Figure 6, we show that the circular velocity of the halo (V, black points) exhibits a turndown in the bTFR at M M⊙, corresponding to velocities 50 km s-1, as galaxy formation becomes inefficient. Our bTFR trend is consistent with the turndown predicted by Sales et al. (2017) for the halo circular speed in the APOSTLE simulations (orange solid line).
@@
We can also calculate the halo’s rotation speed using the midplane gravitational potential (V, black points – top right panel in Figure 6) instead of the enclosed mass. If galaxies are in dynamic equilibrium and the assumption of spherical symmetry is valid, the two methods should return the same velocity. However, as explored in a number of previous works (e.g., Read et al., 2016; El-Badry et al., 2017; Verbeke et al., 2017; Downing & Oman, 2023; Sands et al., 2024), dwarf galaxies have bursty star formation histories with feedback that frequently drives them out of dynamical equilibrium. This effect is likely to become more pronounced as galaxy mass decreases and the gravitational potential well becomes shallower and more susceptible to feedback. Additionally, velocity dispersion becomes more dominant as galaxy mass decreases and rotation velocity declines, such that circular velocity is no longer an accurate measure of velocity. Because of these effects, using the midplane gravitational potential may be a more accurate tracer of the underlying mass distribution, and perhaps more comparable to what an observer might measure at lower rotation speeds. However, for our sample the overall differences in the two methods are subtle (as seen in Figure 2). Differences between the mass enclosed and midplane methods tend to be more significant at smaller radii, but our measurements for the maximum halo speed are measured at larger radii and hence both methods produce similar results for the bTFR. When using V, we again find a turndown at M M⊙.
@@
bTFR using spatially resolved methods: At an H i surface density of 1 M⊙ pc-2, we measure the 2D projected radius, R. Then, we find the maximum rotation speeds for each rotation curve method at radii within R, which we call V (which uses the enclosed mass) and V (which uses the midplane potential). The top left panel of Figure 6 shows the bTFR when we use V (blue triangle markers), while the top right panel shows the bTFR when using V (cyan diamond markers). We compare our spatially resolved results to the V trend in Lelli et al. (2019) (grey dashed line), which uses the maximum velocity from observed rotation curves, for galaxies that have both flat and rising rotation curves. From their Table 1, Lelli et al. (2019) measured a slope = 3.52, intercept = 2.59, and vertical scatter = 0.27 (grey shaded region) with a sample of 153 galaxies. Both spatially resolved H i methods are consistent with observations as they each follow a power-law trend and do not exhibit a turndown. We show representative error bars on the lowest-mass galaxy, in which we estimate the observational errors for velocity and mass (discussed further in Section 3.4).
@@
bTFR with H i Linewidths: In the bottom panel of Figure 6, we show the bTFR when using spatially unresolved H i linewidths. We compare the linewidths (square markers; W in maroon, W in red, and W in salmon) to V (black markers). The observational sample compared here is different than in the other two panels. The dashed line shows the bTFR result from Bradford et al. (2016), which uses the W linewidths for 930 isolated galaxies and has a slope = 3.24, intercept = 3.21, and vertical scatter = 0.25 (grey shaded region). As established by other works (Macciò et al., 2016; Brooks et al., 2017; Dutton et al., 2019; Sardone et al., 2024), W often underestimates the halo rotation speed (V). W and W trace the halo at higher masses, but do not exhibit the bTFR turndown at lower masses. All of our linewidths are from the H i profile when viewing the galaxy edge-on (), and therefore are inclination-corrected.
@@
Errors on the lowest-mass galaxy are based on percent errors in single-dish data, and this is discussed further in Section 3.4. We plot the variance in our simulations on the right side of the plot for W (maroon), W (red), and W (salmon) for each mass bin spanning 1 order of magnitude. H i linewidths vary depending on the viewing angle and mass of the galaxy. We calculate the difference between the linewidth when the galaxy is viewed edge-on versus at a random inclination angle, and then average over all of the galaxies in a given mass bin. Despite the uncertainty on inclination angle for the lower-mass, irregular galaxies, the variance error bars are smallest in this regime as the Gaussian H i profiles do not vary much with viewing angle. In this regime, the variance is small compared to the estimated observational errors.
In summary, a turndown in the bTFR is predicted no matter the method used to find the halo’s maximum rotation speed, either the mass-enclosed (black points, top left panel) or gravitational midplane potential (black points, top right panel). However, the maximum halo value is not generally recoverable with H i , whether for spatially resolved methods at the limit M⊙ pc-2, or the unresolved H i linewidths. For resolved rotation curves, using the enclosed mass to estimate the speed results in only slightly higher rotation speeds for lower-mass galaxies as compared to using the midplane potential (at most, difference), but neither shows a turndown when measuring the maximum resolved H i velocity. For the unresolved H i profiles, no measurement is able to recover the underlying turndown in the bTFR. These results have major implications for observational studies. We further investigate the physical origins of these results, and their implications, below.
3.3 bTFR Slopes and Fit Parameters
@@
For each velocity method, Table 1 shows fits to the equation
@@
| Velocity () | Slope () | Intercept () | |
| V | 0.17 | ||
| V | 0.15 | ||
| W | 0.20 | ||
| W | 0.22 | ||
| W | 0.33 |
For the spatially resolved velocity methods, our bTFR slope using V () is consistent with studies in a similar baryonic mass range. Lelli et al. (2019) found a bTFR slope of , and Šiljeg et al. (2024) found that classical dwarfs observered with Apertif follow this relation. Deg et al. (2024) used early WALLABY observations and found a slope of . Our value is less steep than the slope from Stark et al. (2009) ( (random) (systematic)), and this is likely due to their sample having fewer low-mass galaxies (M M⊙). Stark et al. (2009) also select for gas-rich galaxies that reach a flat rotation curve, which has been shown to produce a slope closer to 4 (Bradford et al., 2016).
For the spatially unresolved methods, our bTFR slope for W () is shallower, while the slopes for W () and W () are closer to observations. Using W, Bradford et al. (2016) measured a slope of , and Lelli et al. (2019) measured . We expect W and W to have steeper bTFR slopes than W, since W < W < W based on the shape of H i profiles. This is consistent with previous studies (Bradford et al., 2016; El-Badry et al., 2017; Lelli et al., 2019; Sardone et al., 2024). Our bTFR slopes from H i linewidths may not exactly match observations due to differences in how we treat inclination angle (as discussed in Section 2.3) and the inclusion of more irregular, lower-mass galaxies in our sample.
We also determine best-fit relationships between the H i velocities and the maximum rotation speeds, V and V. For V, which uses the mass-enclosed gravitational potential, the equations relating the maximum halo speed to the H i velocities are:
The equations relating V (assuming the midplane gravitational potential) to the H i velocities are:
Finally, we determine the relation between the maximum halo speeds V and V:
3.4 Estimate of Observational Errors
The bTFR compares a galaxy’s baryonic mass (MM⋆+1.4M) and rotation velocity, and in this section we elaborate on our errors for each axis. We do not perform mock observations for our simulated dwarf galaxies. However, we attempt to plot our bTFR trends with reasonable errors to determine the significance of differences between H i and halo velocities.
Our baryonic mass error must account for errors in stellar mass and H i mass. For stellar mass, we use an uncertainty of 0.12 dex (33%) adopted by McQuinn et al. (2022). This uncertainty considers the variance in mass-to-light ratios ranging from 0.45-0.6 at 3.6 m, which is the upper-end of uncertainties from calibrations like Zhu et al. (2010), McGaugh & Schombert (2014), and Meidt et al. (2014), resulting in a fractional uncertainty of . For H i mass errors, we adopt the error function from Haynes et al. (2018), which propagates distance () and flux () percent uncertainties:
We use a signal-to-noise ratio of 6.5 ( error), and this is the threshold used by Haynes et al. (2018). We adopt a distance uncertainty of , which is the average for the sample in McQuinn et al. (2022). Distance uncertainties in this range require primary distance methods such as the tip of the red giant branch. Based on this distance uncertainty, we find an H i mass uncertainty of 0.08 dex (). Finally, we take the log-space stellar and H i mass errors in quadrature to get 0.15 dex () as the baryonic mass error.
For the total error in velocity, we consider the uncertainty from turbulent motion (dispersion) and the uncertainty in measuring the rotation speed. The asymmetric drift correction from gas turbulence is usually approximated with a value of km s-1 (Stilp et al., 2013b). By taking the logarithm of the upper limit as a fraction of the fiducial value, we get dex (25% error) as the dispersion uncertainty. For the uncertainty of spatially resolved velocities, we use 0.05 dex (), which is the average for galaxies from McQuinn et al. (2022). We take the log-space errors for dispersion and rotation speed in quadrature to get 0.11 dex (29%) for the velocity error (Figure 6, top panels). For the uncertainty of H i linewidths, we use 0.04 dex (9%), which is the average from Bradford et al. (2016). Adding the errors in quadrature gives a total velocity uncertainty of 0.10 dex (27% error), as shown in the bottom panel of Figure 6.
@@ 3.5 Comparing Velocity Methods and their Radial Extent
We examine the radial extent probed by each method in Figure 7. We plot ratios of the H i velocities and radii relative to the halo values as a function of baryonic mass.
@@
The top row of Figure 7 shows how well the spatially resolved H i methods recover the maximum halo values, whether using the mass enclosed (top left panel) or midplane potential (top right panel). Both methods underpredict the maximum halo speed and corresponding radius by very similar factors at a given baryonic mass. With galaxies at MM⊙, the H i size at the 1 M⊙ pc-2 limit (R) extends to the flat part of the rotation curve. This is apparent with the velocity ratios being close to 1, while the corresponding radius can underestimate the R or R values with factors of . At M M⊙, the H i sizes shrink and limit the radial extent of spatially resolved rotation curves. R moves interior to the rising part of the rotation curve, and the H i velocities significantly underpredict the halo speed. The velocity differences can be close to 50% for both the mass-enclosed and midplane methods.
The bottom right panel in Figure 7 demonstrates how the midplane and mass-enclosed H i velocities differ with baryonic mass. As stated earlier, if galaxies are in dynamic equilibrium and the assumption of spherical symmetry holds, the two methods should return the same velocity, but we find slight differences. At M, V systematically measures lower velocities than V with up to a difference, while at higher masses, V measures higher velocities with up to a 6.2% difference. This trend is consistent with the dynamical expectation that a disky mass distribution, as found in the more massive dwarfs, will peak at a slightly higher velocity than a spherical distribution. Our the lower-mass galaxies tend to be more prolate (see Keith et al., 2025), and will be more prone to disequilibrium due to feedback given their shallower potential wells. These mass trends explain the velocity trend seen in the lower right panel.
In the bottom left panel of Figure 7, we compare the velocity and radius ratios for W/2. Technically, the unresolved H i linewidths are not set to a size extent of the measured HI, as the entire dwarf is usually detected within the single dish beam size. However, we define the corresponding (mass-enclosed) circular velocity as that which gives the same velocity as W/2. We find that this measures an equivalent radius that is comparable to the sizes from resolved rotation curves. Compared to the other H i linewidths, W is more sensitive to diffuse H i gas at larger radii and rotation speeds (Sardone et al., 2024). Given the Gaussian shape of the H i profile in dwarf galaxies, Sardone et al. (2024) found that W provides a larger velocity, closer to the maximum halo speed, though likely still an underestimate. In observations, W can be susceptible to noise and requires very deep and high signal-to-noise measurements. Here, we show our (noise-free) predictions for how W relates to V. We find that W differs up to %. Compared to V and VW has more vertical scatter in its velocity ratios. Since we plot all H i linewidths with , this vertical scatter is likely not due to inclination, but it might be from diversity in H i profile shapes. Turbulent velocities have a non-negligible impact on H i linewidth shape, especially at lower galaxy masses (e.g., El-Badry et al., 2018). We will test the impact of dispersion on H i linewidths in future work.
In all methods, the velocity ratios and radius ratios decrease with lower baryonic masses. As galaxy mass decreases, H i velocities more significantly underestimate the maximum halo speed. A lower rotation speed is due to the fact that the H i radius shrinks and traces a smaller radius on the rotation curve, on the rising part. By comparing the velocity and radius ratios for the lowest mass galaxy in each panel of Figure 7, we can infer limiting factors for each method. For V, we assume the same (mass-enclosed) gravitational potential for the H i rotation curve and halo speed, and therefore the main limiting factor must be the observable extent of the H i disk. The midplane method (V) measures at the same radius, but gives a slightly lower velocity than V as mass decreases. This suggests that in addition to the H i size, V measurements also differ from V due to non-circular motions and dynamical disequilibrium. Lastly, as discussed above, W seems to be the most inaccurate tracer of the three presented in Figure 7, and it is not immediately clear why. We will explore the physical processes that set H i line profile shape in a future work.
@@
Previously, we determined that our galaxies with MM⊙ have canonical H i disk sizes which have reached the flat part of the rotation curve. However, the H i disk size is sensitive to the surface density limit that can be achieved by observations (e.g., Wang et al., 2025). So far we have used the H i limit of M⊙ pc-2, which is the standard definition used in the literature. New and future H i surveys can push to deeper limits (as discussed in Section 4.2). In this section, we explore the sensitivities required to trace R and subsequently V.
@@
Figure 8 shows the H i column density profiles in three stellar mass bins for our simulated dwarf galaxies. Horizontal lines correspond to the H i surface density limits: 1.0, 0.08, 0.02, and 0.01M⊙pc-2. Our column density profiles plateau with radius, but with different plateaus as a function of mass. More massive galaxies (MM⊙) plateau around cm-2, and less massive galaxies plateau around cm-2. Our sample exhibits the full range of H i column densities found in other simulations ( cm-2), in which the lower-density gas corresponds to cold accretion from the intergalactic medium (Kereš et al., 2005; Popping et al., 2009; Cook et al., 2024).
@@
In Figure 9, we compare our surface density profiles to the median surface density profile for the FEASTS sample (Wang et al., 2025). This profile is normalized by R, which is the H i size at a surface density limit of 0.01 M⊙ pc-2. The FEASTS sample targets more massive galaxies, with M M⊙ and M M⊙, while most of our galaxies are below this mass range. Thus, it is not clear if this comparison is valid, but the normalized R/R span two orders of magnitude in galaxy mass and show uniformity, suggesting extrapolation to lower masses might be acceptable. With this caveat, we find that the observed median profile lies within the range of our simulated surface density profiles. It seems that the simulations might have larger H i extents compared to the observed sample, though this interpretation appears to be at odds with the results shown in Figure 4. However, discrepancies may also be due to the feedback and radiative transfer models used in our simulations. For example, simulations have a hard time maintaining molecular gas, which wants to immediately form star particles, so must be regulated by feedback. Our higher H i surface densities in the central regions of galaxies are likely because some of this gas should be molecular, but feedback turns it to H i instead. Overall, it is clear that the simulations tend to follow the observational trend in normalized H i surface density profile at large radii. Additional observations of surface density profiles out to large radii like R will help us to calibrate the gas physics.
@@
In Figure 10, we compare the bTFR using V at different surface density limits. As expected, for galaxies with V km s-1, the higher surface density of 2 M⊙ pc-2 (equivalent to worse sensitivity) probes a radial extent that is interior to the radius corresponding to 1 M⊙pc-2 (see Figure 8). Thus, the measured H i disk radius and rotation velocity both get smaller, because the velocity measurement moves inward on the rising part of the rotation curve. On the other hand, improving sensitivity and reaching a lower H i column density does indeed begin to close the gap between V and V. At 0.08 M⊙ pc-2, V recovers V down to 45 km s-1. This leads to a change in the bTFR slope as a function of H i surface density limit, which can be tested by observations. We fit one power-law slope for our sample of galaxies (with V km s-1). For 0.08 M⊙ pc-2, the bTFR slope becomes steeper with a value of , and steeper still for 0.01 M⊙ pc-2 with a value of . As demonstrated by the lowest limits, a slope may indicate a turndown in the bTFR. Measuring the outermost rotation velocity at lower surface densities could allow us to recover the maximum halo speed, though sufficient spatial resolution is also needed. At a surface density limit of 0.01 M⊙ pc-2, our sample has H i sizes ranging from kpc.

### v2

@@ 2.2 Spatially unresolved H i profiles
We use a fixed velocity resolution of km s-1 and measure each galaxy over the velocity range -200 to 200 km s-1. This channel width is set to match the best spectral resolution in Bradford et al. (2015). Our data cubes have fixed spatial axes spanning from -100 to 100 kpc across each galaxy with a resolution of 3.7 kpc, though the spatial resolution has little impact on the unresolved linewidth results. We do not consider beam smearing effects or observational noise, but will include these effects in a future work. The emission spectra are converted from H i mass to flux by inverting the relation
where is the velocity in km s-1, is the spectrum in Jy beam-1, and we set the distance Mpc. The H i profiles of our full sample are shown in Appendix A.
@@ 2.3 Inclinations
At M M⊙, our simulated dwarfs are predominantly irregular rather than disky (Keith et al., 2025). For these lowest-mass dwarf galaxies, we have found that our orientation method does not always yield a result that is edge-on by visual inspection. It is possible that there is no orientation that yields an apparent edge-on result, given their irregular nature. Although the inclinations for these irregular galaxies are difficult to measure and correct for, the Gaussian H i profiles of these lower-mass dwarf galaxies do not vary much by viewing angle (as quantified by the errors to the right of each panel in the bottom row of Figure 6. Our inclination procedure with simulations allows us to include dwarf galaxies that would otherwise be removed from observational samples.
@@ 2.4 Definitions for Velocities and Radii
There are a number of different ways to measure rotation velocity, and at different radii. The methods for velocities and corresponding radii in this work are:
@@
W and R: W, W, and W are the H i profile’s width at different percentages of the peak flux height; 10%, 20%, and 50%, respectively. Technically, the unresolved H i linewidths are not set to a size extent, as the entire dwarf is usually detected within the single dish beam size. However, we calculate an equivalent radius by using the corresponding radius (R) where the linewidth W intersects the circular rotation curve. In some cases, W does not intersect the rotation curve (see Figure 2, top panel as an example). H i linewidths may not exactly correspond to the circular velocity if the emission profile is asymmetric. If the gas is more extended on one side of the galaxy, hence moving at a faster speed, this results in a higher linewidth. Therefore, this linewidth can exceed the max velocity when assuming spherical shells of the same mass enclosed.
@@
In this work, we aim to understand how the different methods and velocity definitions impact the determined bTFR. We do not perform full mock observations, i.e., we do not include noise, beam smearing, or other observational effects in our H i data cubes. We also do not take into account the impact of non-circular motions, though we discuss the implications of this in Section 4.2. We assess whether the various velocity measurement methods have the potential to recover a turndown in the bTFR even with these ‘idealized’ rotation curves and H i linewidths.
@@ 3.1 Comparing Stellar and H i Properties to Observations
After confirming that our masses are consistent with observations, we check the H i size-mass relationship in the bottom panel of Figure 4. The size is defined as the diameter where the H i surface density has dropped to 1 M⊙ pc-2, similar to observations. We compare the H i size-mass relation from the simulated dwarf galaxies to the observations from Wang et al. (2016), which follows the trend from Broeils & Rhee (1997) (B97), and has been confirmed by recent WALLABY (Deg et al., 2024) and FEASTS results (Wang et al., 2025). Our galaxies fit in the 3- scatter of Wang et al. (2016), represented by the solid black lines. Stevens et al. (2019) found that the H i size-mass relation is robust to gas stripping and galaxy population (centrals versus satellites), so matching this relation mostly confirms that our feedback is of appropriate strength and does not disrupt the H i disk.
@@
bTFR using the maximum halo speed: Modeling the halo’s circular velocity based on the total mass enclosed in spherical shells is common in bTFR studies (e.g., Trujillo-Gomez et al., 2018; McQuinn et al., 2022; Lelli et al., 2019), so we use this definition as our reference point. In Figure 6, we show that the circular velocity of the halo (V, black points) exhibits a turndown in the bTFR at M M⊙, corresponding to velocities 50 km s-1, as galaxy formation becomes inefficient. Our bTFR trend is consistent with the turndown predicted by Sales et al. (2017) for the halo circular speed in the APOSTLE simulations (orange solid line).
@@
We can also calculate the halo’s rotation speed using the midplane gravitational potential (V, black points – top right panel in Figure 6) instead of the enclosed mass. If galaxies are in dynamic equilibrium and the assumption of spherical symmetry is valid, the two methods should return the same velocity. However, as explored in a number of previous works (e.g., Read et al., 2016; El-Badry et al., 2017; Verbeke et al., 2017; Downing & Oman, 2023; Sands et al., 2024), dwarf galaxies have bursty star formation histories with feedback that frequently drives them out of dynamical equilibrium. This effect is likely to become more pronounced as galaxy mass decreases and the gravitational potential well becomes shallower and more susceptible to feedback. Additionally, velocity dispersion becomes more dominant as galaxy mass decreases and rotation velocity declines, such that circular velocity is no longer an accurate measure of velocity. Because of these effects, using the midplane gravitational potential may be a more accurate tracer of the underlying mass distribution, and perhaps more comparable to what an observer might measure at lower rotation speeds. However, for our sample the overall differences in the two methods are subtle (as seen in Figure 2). We further discuss the effect of non-circular motions and the bTFR in Section 4.2. Differences between the mass enclosed and midplane methods tend to be more significant at smaller radii, but our measurements for the maximum halo speed are measured at larger radii and hence both methods produce similar results for the bTFR. When using V, we again find a turndown at M M⊙.
As mentioned in Section 2, the superbubble feedback used in the Massive Dwarfs simulations (black circles) may result in slightly higher halo masses. A close examination of the black points in Figure 6 suggests that the higher halo masses lead to slightly higher V values in the Massive Dwarfs. In other words, the use of the combined simulations may broaden the scatter in the simulated bTFR. Despite this, the variance in max halo speeds between feedback models is small, such that the spread introduced is still within the observational error. This is most obvious when comparing the V spread in the H i results (colored points, discussed in detail below) relative to the spread in observations (grey shaded regions). The change due to feedback is small enough to not affect the conclusions of this work, and considering the effects of feedback on the bTFR turndown is beyond the scope of this paper.
@@
bTFR using spatially resolved methods: At an H i surface density of 1 M⊙ pc-2, we measure the 2D projected radius, R. Then, we find the maximum rotation speeds for each rotation curve method at radii within R, which we call V (which uses the enclosed mass) and V (which uses the midplane potential). The top left panel of Figure 6 shows the bTFR when we use V (blue markers), while the top right panel shows the bTFR when using V (cyan markers). We compare our spatially resolved results to the V trend in Lelli et al. (2019) (grey dashed line), which uses the maximum velocity from observed rotation curves, for galaxies that have both flat and rising rotation curves. From their Table 1, Lelli et al. (2019) measured a slope = 3.52, intercept = 2.59, and vertical scatter = 0.27 (grey shaded region) with a sample of 153 galaxies. Both spatially resolved H i methods are consistent with observations as they each follow a power-law trend and do not exhibit a turndown. We show representative error bars on the lowest-mass galaxy, in which we estimate the observational errors for velocity and mass (discussed further in Appendix B).
@@
bTFR with H i Linewidths: In the bottom row of Figure 6, we show the bTFR when using spatially unresolved H i linewidths. We compare the linewidths (bottom left panel: W in red markers, bottom right panel: W in maroon markers) to V (black markers). The observational sample compared here is different than in the previous two panels. The dashed line shows the bTFR result from Bradford et al. (2016), which uses the W linewidths for 930 isolated galaxies and has a slope = 3.24, intercept = 3.21, and vertical scatter = 0.25 (grey shaded region). As established by other works (Macciò et al., 2016; Brooks et al., 2017; Dutton et al., 2019; Sardone et al., 2024), W often underestimates the halo rotation speed (V). W and W trace the halo at higher masses, but do not exhibit the bTFR turndown at lower masses. All of our linewidths are from the H i profile when viewing the galaxy edge-on (), and therefore are inclination-corrected.
@@
Errors on the lowest-mass galaxy are based on percent errors in single-dish data, and this is discussed further in Appendix B. We show the variance in our simulations on the right side of the plot for W and W for each mass bin spanning 1 order of magnitude. H i linewidths vary depending on the viewing angle and mass of the galaxy. We calculate the difference between the linewidth when the galaxy is viewed edge-on versus at a random inclination angle, and then average over all of the galaxies in a given mass bin. Despite the uncertainty on inclination angle for the lower-mass, irregular galaxies, the variance error bars are smallest in this regime as the Gaussian H i profiles do not vary much with viewing angle. In this regime, the variance is small compared to the estimated observational errors.
@@
For each velocity method, the information under the legends in Figure 6 shows fits to the equation
@@
The fits are shown as dotted colored lines in each panel of Figure 6.
In summary, a turndown in the bTFR is predicted no matter the method used to find the halo’s maximum rotation speed, either the mass-enclosed (black points, top left panel) or gravitational midplane potential (black points, top right panel). However, the maximum halo value is not generally recoverable with H i , whether for spatially resolved methods at the limit M⊙ pc-2, or the unresolved H i linewidths. For resolved rotation curves, using the enclosed mass to estimate the speed results in only slightly higher rotation speeds for lower-mass galaxies as compared to using the midplane potential (at most, difference), but neither shows a turndown when measuring the maximum resolved H i velocity. For the unresolved H i profiles, no measurement is able to recover the underlying turndown in the bTFR. These results have major implications for observational studies, as discussed in Section 4.3.
@@ 3.3 Comparing Velocity Methods and their Radial Extent
In Figure 7, we examine the radial extent probed by each method. We plot ratios of the H i velocities and radii relative to the halo values as a function of baryonic mass. Both spatially resolved methods (left and middle columns in Figure 7) underpredict the maximum halo speed and corresponding radius by very similar factors at a given baryonic mass. With galaxies at M M⊙, the H i size at the 1 M⊙ pc-2 limit (R) extends to the ‘flat’ part of the rotation curve. This is apparent with the velocity ratios being close to 1 above this mass. At lower masses, the H i size shrinks as galaxy mass decreases, and limits the radial extent of spatially resolved rotation curves. As the H i shrinks, it no longer extends to the radius where the halo’s maximum velocity is reached. This shrinking of H i size is the primary reason that the spatially resolved H i methods under-predict the maximum rotation speed of the lowest-mass galaxies.
In the right column of Figure 7, we compare the max velocity and corresponding radius of W/2 to the halo value. We measure an equivalent radius (R) corresponding to the (mass-enclosed) circular velocity curve at a value equal to W/2. Compared to V and V, W has more vertical scatter in its velocity and radius ratios. Since we plot all H i linewidths at , this vertical scatter is likely not due to inclination, but perhaps from diversity in H i profile shapes (see Figure 11). We will study the impact of dispersion on H i linewidths in future work.
In Appendix C, we also determine best-fit relationships between the H i velocities and the maximum rotation speeds, V and V, which could potentially be used to extrapolate observational results.
@@
We determined that our galaxies with MM⊙ have canonical H i disk sizes which have reached the flat part of the rotation curve. However, the H i disk size is sensitive to the surface density limit in observations (e.g., Wang et al., 2025). So far we have used an H i limit of M⊙ pc-2, which is the standard definition used in the literature. New and future H i surveys can push to deeper limits (as discussed in Section 4.3). In this section, we explore the sensitivities required to trace R and subsequently V.
@@
Figure 8 shows the H i column density profiles in three stellar mass bins for our simulated dwarf galaxies. Horizontal lines correspond to the H i surface density limits: 1.0, 0.08, 0.02, and 0.01M⊙pc-2. Our column density profiles plateau with radius, but with different plateaus as a function of mass. More massive galaxies (MM⊙) plateau around cm-2, and less massive galaxies plateau around cm-2. These column density profiles are calculated with Pynbody using radial shells. Using the same simulations, Piacitelli et al. (2025) calculated column density profiles with a different method, using sightlines throughout the circumgalactic medium. We have verified that we are consistent out to the 20 kpc shown here. Our simulated galaxies exhibit the full range of H i column densities found in other simulations ( cm-2), in which the lower-density gas corresponds to cold accretion from the intergalactic medium (Kereš et al., 2005; Popping et al., 2009; Cook et al., 2024).
@@
In Figure 9, we compare our surface density profiles to the median surface density profile for the FEASTS sample (Wang et al., 2025). This profile is normalized by R, which is the H i size at a surface density limit of 0.01 M⊙ pc-2. The FEASTS sample targets more massive galaxies, with M M⊙ and M M⊙, while most of our galaxies are below this mass range. Thus, it is not clear if this comparison is valid, but the normalized R/R span two orders of magnitude in galaxy mass and show uniformity, suggesting extrapolation to lower masses might be acceptable. With this caveat, we find that the observed median profile lies within the range of our simulated surface density profiles. Galaxies from the Massive Dwarfs simulations appear to have excess surface density of H i in their inner regions compared to observations. Discrepancies may be due to the feedback and radiative transfer models used in our simulations. For example, simulations have a hard time maintaining molecular gas, which wants to immediately form star particles, so must be regulated by feedback. Our higher H i surface densities in the central regions of galaxies are likely because some of this gas should be molecular, but feedback turns it to H i instead. Overall, it is clear that the simulations tend to follow the observational trend in normalized H i surface density profile at large radii. Additional observations of surface density profiles out to large radii like R will help us to calibrate the gas physics.
@@
In Figure 10, we compare the bTFR using V at different surface density limits. Improving the sensitivity and reaching a lower H i column density does indeed begin to close the gap between V and V. At 0.08 M⊙ pc-2, V recovers V down to 45 km s-1. This leads to a change in the bTFR slope as a function of H i surface density limit, which can be tested by observations. We fit one power-law slope for our sample of galaxies (with V km s-1). For 0.08 M⊙ pc-2, the bTFR slope becomes steeper with a value of , and steeper still for 0.01 M⊙ pc-2 with a value of . As demonstrated by the lowest limits, a slope may indicate a turndown in the bTFR. The bTFR slope may be more indicative of our sensitivity limits than alternate models of gravity, as we can measure a slope within CDM. Measuring the outermost rotation velocity at lower surface densities allows us to recover the maximum halo speed, though sufficient spatial resolution is also needed. At a surface density limit of 0.01 M⊙ pc-2, our sample has H i sizes ranging from kpc.